# Ultrafast Laser-Fabricated Fluoride Glass Waveguides with Exceptionally High Refractive Index Change for Mid-Infrared Integrated Optics

### v1

@@ 4 Conclusion
In conclusion, this study marks a significant milestone in the development of integrated optical elements in fluoride glasses for both visible and mid-infrared applications. Overcoming decades of challenges, the research has yielded optical waveguides with an unprecedented positive refractive index change, exceeding . The investigation into the origins of the index change is attributed to material densification driven mainly by the migration of barium in the revised fluoride glass composition. Such a strong migration to form optical waveguides eluded in the past when waveguides were inscribed in the general composition which contained lanthanum fluoride. The current recomposition ensured the migration of heavy barium atoms to positive index zone densifying with an increase in local index. The optimal fabrication conditions involved energy per pulse within the range of 2 to 6 J, feedrates spanning from 40 to 150 m/s, and a multiscan pitch of 1 m. The waveguides were approximately 12 m wide and 14 m tall, which then yielded a multimoded behaviour at 3.5 m wavelength. The losses were as low as 0.21 dB/cm which were measured at a wavelength of 1.62 m. A broad band butt coupled HR mirror and a fiber bragg grating was used to form a cavity to lase these active waveguides at 1542.83 nm wavelength.

### v2

@@ 4 Conclusion
In conclusion, this study marks a significant milestone in the development of integrated optical elements in fluoride glasses for both visible and mid-infrared applications. Overcoming decades of challenges, the research has yielded optical waveguides with an unprecedented positive refractive index change, exceeding . The investigation into the origins of the index change is attributed to material densification driven mainly by the migration of barium in the revised fluoride glass composition. Such a strong migration to form optical waveguides eluded in the past when waveguides were inscribed in the composition which contained lanthanum fluoride. The current recomposition ensured the migration of heavy barium atoms to positive index zone densifying with an increase in local index. The optimal fabrication conditions involved energy per pulse within the range of 2 to 6 µJ, feedrates spanning from 40 to 150 µm/s, and a multiscan pitch of 1 µm. The waveguides were approximately 12 µm wide and 14 µm tall, which then yielded a multimoded behaviour at 3.5 µm wavelength. The losses were as low as 0.21 dB/cm which were measured at a wavelength of 1.62 µm. A broad band butt coupled HR mirror and a fiber bragg grating was used to form a cavity to lase these active waveguides at 1542.83 nm wavelength.